OpenStudy (anonymous):

Please help, will give medal! I just need someone to check my work. I know what the answer should be but it keeps coming out wrong. In the problems prior to this one, we are given some of the information that I've already plugged into the formula. Basically I'm finding the volume of a sphere and then dividing it by four.

OpenStudy (anonymous):

5. The theme park company is building a scale model of the killer whale stadium main show tank for an investor's presentation. Each dimension will be made 6 times smaller to accommodate the mock-up in the presentation room. How many times smaller than the actual volume is the volume of the mock-up? You can find the answer by dividing the old volume by the new volume. So first we find the new volume of the tank by taking its dimensions (it’s radius) and dividing it by six. new radius = (old radius)/6 new radius = 70/6 new radius = 11.6666666667 And with the new radius, we can find the volume of the entire sphere… V = 4/3πr^2 V = 4/3π11.6666666667^2 V = 4/3π136.111111112 V = 4/3 • 427.605666741 The answer should be V = 6651.64 but for some reason my calculator is giving me something like 570 while everything else is giving me 6651

jimthompson5910 (jim_thompson5910):

The volume of a sphere is \[\Large V = \frac{4}{3}\pi r^3\] notice how the exponent over 'r' is 3 and NOT 2

OpenStudy (anonymous):

Oh my gosh, I made a careless mistake. Thank you! X) I feel dumb.

jimthompson5910 (jim_thompson5910):

no worries